OpenStudy (anonymous):

y=3x-x2= I am try to find the intercepts and then sketch this problem

OpenStudy (anonymous):

Just FYI, \[x ^{2}\] can be written x^2 online (which is what I'm assuming you meant above). 1. Factor the 3x-x^2 so you have y=x(3-x). 2. To find y-intercepts, set x=0. When x=0, you have y=0*3=0. 3. To find x-intercepts, set y=0 and solve for x. By the zero-product rule, y=0 when x=0 or x-3=0. Therefore, you have x-intercepts at x=0 and x=3.